OpenStudy (anonymous):

Jane walked 4/5 mile on Monday. On Teusday, she walked1/2 of the distance she walked on Monday. How many miles did jane walk on Teusday. A. 2/5 mile B. 1/5 mile C.1/2 mile D. 3/4 mile

OpenStudy (anonymous):

I think it is A. but i am not sure

OpenStudy (mystiz):

it is A

OpenStudy (mystiz):

look

OpenStudy (mystiz):

4/5 x 1/2 = 4/10 = 2/5

OpenStudy (anonymous):

Thank you

OpenStudy (mystiz):

no problem